# Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages

### v1

@@ Abstract
Quantifying reasoning capability in low-resource languages remains a challenge in NLP due to data scarcity and limited access to annotators. While LLM-assisted dataset construction has proven useful for medium- and high-resource languages, its effectiveness in low-resource languages, particularly for commonsense reasoning, is still unclear. In this paper, we compare three dataset creation strategies: (1) LLM-assisted dataset generation, (2) machine translation, and (3) human-written data by native speakers, to build a culturally nuanced story comprehension dataset. We focus on Javanese and Sundanese, two major local languages in Indonesia, and evaluate the effectiveness of open-weight and closed-weight LLMs in assisting dataset creation through extensive manual validation. To assess the utility of synthetic data, we fine-tune language models on classification and generation tasks using this data and evaluate performance on a human-written test set. Our findings indicate that LLM-assisted data creation outperforms machine translation.
Synthetic Data Generation for Culturally Nuanced Commonsense Reasoning in Low-Resource Languages
@@
Figure 1: Proposed pipeline, including the train set overgeneration process
@@ 1 Introduction
Reasoning—the ability to draw conclusions, make inferences, and understand relationships between concepts based on available information—is a key capability assessed in recent developments of large language models (Dubey et al., 2024; OpenAI et al., 2024; Hurst et al., 2024; Almazrouei et al., 2023). It has been widely studied in English through various benchmarks, including StoryCloze Mostafazadeh et al. (2016, 2017), WinoGrande Sakaguchi et al. (2021), and HellaSwag Zellers et al. (2019), which evaluate different aspects of reasoning such as story comprehension, pronoun resolution, and natural language inference. Beyond English, research on reasoning has largely relied on machine translation Ponti et al. (2020); Lin et al. (2022), often overlooking the cultural context of the target languages. While some datasets have been developed from scratch for medium-resource languages like Indonesian Koto et al. (2024), reasoning in low-resource languages remains significantly underexplored. The primary challenges include the high cost of annotation, difficulties in recruiting native speakers, and the limited availability of culturally relevant text.
In this paper, we ask: to what extent can large language models (LLMs) assist in creating reasoning datasets for low-resource languages? How do they compare to machine translation and building datasets from scratch? While LLM-assisted dataset construction has been widely explored, most efforts have focused on English Liu et al. (2022, 2024a) and traditional NLP tasks such as named entity recognition (Mohit, 2014) and question answering Ghazaryan et al. (2025); Putri et al. (2024). However, their potential for generating reasoning datasets in low-resource languages remains under-studied, particularly for culturally nuanced story comprehension tasks.
Unlike sentence-pair classification tasks for assessing reasoning (e.g., the COPA dataset), we take a story comprehension approach to evaluate LLMs’ ability to generate coherent narratives in Javanese and Sundanese over longer texts (five sentences). Specifically, we follow the StoryCloze framework Mostafazadeh et al. (2016, 2017) to assess commonsense reasoning in daily life scenarios, where a four-sentence premise is provided along with two possible options for the concluding sentence (see Figure 1).
We focus on Javanese and Sundanese because, despite having over 80 million and 32 million speakers, respectively Aji et al. (2022), these languages remain underrepresented in language technology development. Beyond their linguistic diversity, both languages are deeply intertwined with distinct cultural traditions that influence their usage. Sundanese, for instance, conveys social nuanced through varying degrees of politeness and phonetic distinctions, while Javanese employs a structuted system of speech levels to reflect hierarchical relationships in communication (Wolff and Poedjosoedarmo, 1982).
Our contribution can be summarized as follows:
  • We benchmark three open-weight and three closed-weight LLMs for StoryCloze-style synthetic data generation in Javanese and Sundanese through extensive human evaluation.
  • We introduce the first commonsense reasoning datasets for Javanese and Sundanese. For training data, we use two strategies: (1) LLM-generated data with few-shot prompting and (2) machine translation. For the test set, we employ two approaches: (1) machine translation with post-editing and (2) manually written samples by native speakers based on predefined topics, with extensive validation to ensure quality. In total, we release 1.22K high-quality LLM-generated samples, 1K human-reviewed machine-translated texts, and 1.12K human-written samples [^1] Our data are available at https://huggingface.co/datasets/rifoag/javanese_sundanese_story_cloze.
  • We conducted extensive experiments in both classification and generation settings to assess the utility of LLM-generated data and machine translation.
Javanese Sundanese
Original
Premise: Ing Yogyakarta, gudeg dadi panganan khas. Gudeg dimasak suwene jam-jaman nganggo nangka enom lan santan. Biasane disajikake karo krecek, endog, lan ayam suwir. Wong Yogyakarta biasane mangan gudeg ing esuk utawa nalika ana acara khusus. Correct: Gudeg diwarisake turun-temurun lan dadi bagian saka budaya kuliner Yogyakarta. Incorrect: Gudeg biasane digoreng nganggo lenga panas lan disajikake pedhes. Premise: Mang Tarya indit ka sawah keur ngarambét. Geus loba jukut nu mucunghul ngaganggu paré. ngarambét ngabaladah biasana dilaksanakeun tilu puluh poé saprak tandur. Mun ngarambét mindo dilaksanakeun lima puluh poe saprak tandur. Correct: Ngarambét dilaksanakeun pikeun ngaberesihan jukut di petakan sawah. Incorrect: Ngarambét dilaksanakeun saméméh panén.
English translation
Premise: In Yogyakarta, gudeg is a traditional dish. It is cooked for hours using young jackfruit and coconut milk. It is usually served with krecek, eggs, and shredded chicken. People in Yogyakarta typically eat gudeg in the morning or during special occasions. Correct: Gudeg is passed down through generations and has become a part of Yogyakarta’s culinary heritage. Incorrect: Gudeg is usually deep-fried in hot oil and served spicy. Premise: Mang Tarya went to the rice field to do ngarambét. There are already lots of wild grass that may harm the rice crops. Ngarambét ngabaladah is done thirty days after tandur. Meanwhile, Ngarambét mindo is done fifty days after tandur. Correct: Ngarambét is done to clean up wild grass in the rice fields. Incorrect: Ngarambét is done before harvest.
Table 1: Examples of story cloze in Javanese and Sundanese, with English translations for illustrative purposes.
@@ 2 Related Works
LLM-Generated Data Creation
One possible solution to tackle data scarcity in NLP is applying data augmentation (Feng et al., 2021; Ding et al., 2020; Ahmed and Buys, 2024; Liu et al., 2024b; Yong et al., 2024; Guo and Chen, 2024; Liu et al., 2022). These methods leverage the generative capabilities of LLMs to produce high-quality synthetic data that can sometimes complement or even substitute manual dataset creation.
Prior studies have explored techniques for enhancing synthetic data generation, primarily for classification tasks. WANLI (Liu et al., 2022) used GPT-3 (Brown et al., 2020) to generate synthetic natural language inference data in English Bowman et al. (2015), followed by human refinement. Yong et al. (2024) leveraged LLMs to generate synthetic English data for sentiment analysis and topic classification, then applied a bilingual lexicon for word-to-word translation into low-resource languages. In contemporaneous work, Putri et al. (2024) utilized GPT-4 OpenAI et al. (2024) for question-answering tasks in one low-resource language. Unlike these works, we focus specifically on reasoning tasks in two low-resource languages and compare multiple strategies, including LLM-assisted generation using both open-weight and closed-weight models, machine translation, and human-written data by native speakers. To ensure topic coverage, we provide predefined topics and example seeds for both LLM-assisted generation and human writing.
English Commonsense Reasoning
Story comprehension has long been a significant challenge in NLP due to the complex interplay of causal and temporal relationships within narratives. A landmark contribution to the evaluation of these aspects is the Story Cloze test, introduced by Mostafazadeh et al. (2016, 2017). Another set of benchmarks for assessing commonsense reasoning includes WinoGrande Sakaguchi et al. (2021), COPA (Gordon et al., 2012), and HellaSwag (Zellers et al., 2019).
WinoGrande (Sakaguchi et al., 2021) is designed to measure the resolution of pronouns in ambiguous contexts. The Choice of Plausible Alternatives (COPA) (Gordon et al., 2012) evaluates causal reasoning by presenting two possible continuations of a given premise, requiring models to infer the most plausible outcome. HellaSwag (Zellers et al., 2019) builds on the SWAG Zellers et al. (2018) dataset by increasing difficulty through adversarial filtering, which ensures that the correct continuation of a given context is not easily identified by simple statistical cues, but rather requires deep reasoning. These benchmarks collectively contribute to a more comprehensive evaluation of commonsense understanding in NLP models.
Commonsense Reasoning in Languages Beyond English
Evaluating commonsense reasoning across multiple languages has gained attention through multilingual benchmarks such as XCOPA (Ponti et al., 2020) and XStoryCloze (Lin et al., 2022). These benchmarks extend existing commonsense reasoning datasets to a wider set of languages by translating English corpora into various target languages. However, such translation-based approaches often fail to capture cultural nuances and may introduce artifacts that affect naturalness and contextual relevance. Additionally, neither XCOPA nor XStoryCloze include benchmarks for Sundanese or Javanese, leaving a gap in evaluating commonsense reasoning in these widely spoken Indonesian local languages.
Several high-quality Indonesian commonsense reasoning datasets have been developed, primarily through human annotation, including IndoCloze Koto et al. (2022a) for narrative understanding, COPAL-ID Wibowo et al. (2024) integrating Indonesian cultural knowledge (i.e., Jakarta region), and IndoCulture Koto et al. (2024) for cultural reasoning. While these resources support commonsense reasoning in Indonesian, a medium-resource language, extensive datasets for local languages remain scarce. In contrast, our work focuses on low-resource languages, specifically Javanese and Sundanese, two of Indonesia’s most widely spoken local languages. We build on IndoCloze as a reference for synthetic data generation and complement it with manually written stories. Unlike prior datasets that rely on MCQs or sentence-pair tasks, we emphasize full story comprehension. To ensure diversity, we adopt multiple data creation approaches, including LLM-generated, machine-translated, and human-written narratives, enabling a broader and more culturally rich evaluation setting.
@@ 3 Dataset Construction
As shown in Figure 1, our dataset construction is describued in two streams: training and testing—using IndoCloze Koto et al. (2022a) as an initial source. IndoCloze is the only story comprehension dataset for commonsense reasoning evaluation in Indonesian and incorporates Indonesian cultural elements. Each sample consists of a four-sentence premise followed by two candidate endings: one correct and one incorrect, representing the fifth sentence.
We use Indonesian dataset in our data construction as it is the national language and shares cultural elements with Javanese and Sundanese, making it a relevant source of seed data. In addition to leveraging IndoCloze, we also create a dataset from scratch by asking native speakers to write and validate samples to ensure cultural and linguistic authenticity. Similar to Mostafazadeh et al. (2016) and Koto et al. (2022a), our dataset focuses on everyday events within Javanese and Sundanese cultural contexts, incorporating relevant locations, names, foods, and cultural practices.
@@ 3.1 Training Data
Our training set is constructed using two strategies: (1) LLM-assisted data generation and (2) machine translation.
LLM-assisted Data Generation
To assess the capability of different LLMs in generating meaningful training data, we create synthetic datasets using three open-weight models: Mixtral-8x7B-Instruct (Jiang et al., 2024), Gemma2-27B-it (Rivière et al., 2024), and Llama3.1-70B (Dubey et al., 2024)—and three closed-weight models: GPT-4o (Hurst et al., 2024; OpenAI et al., 2024), Cohere Command-R-Plus (Cohere, 2024), and Claude-3-Opus (Anthropic, 2024).
For each language and LLM, we provide seed examples and predefined topics. To create seed samples, we manually translate 50 samples from IndoCloze’s training set into Javanese and Sundanese, ensuring cultural localization. For example, we adapt references to food, such as Gudeg (a traditional breakfast from Yogyakarta), and traditions like Sawéran (a ritual in Sundanese wedding ceremonies).
To ensure that the generated examples align with the intended cultural context, we use cultural topics derived from IndoCulture (Koto et al., 2024), covering areas such as food, weddings, family relationships, pregnancy and children, death, religious holidays, agriculture, fisheries and trade, art, traditional games, daily activities, and socio-religious aspects of life. The full prompt used in this study is provided in Figure 10.
Each LLM independently generate 2,000 samples (1,000 per language), resulting in six distinct datasets, each treated as a separate training set. For each LLM call, we randomly select five seed samples and prompted the model to generate new examples. The temperature is set to 0.7 to introduce variability in the generated outputs. The generation process is repeated until each LLM produced at least 1,000 valid examples. On average, each model generated approximately 166 samples per topic (see Appendix D for topic distribution).
Since not all LLM-generated samples are high quality, we train an XLM-R binary classifier to distinguish between good and bad samples. To create training data, we ask workers to rate the quality of 600 LLM-generated outputs (details provided in Section 4.2). These rated samples are then used to train the classifier. [^2] The classifier achieves 68.33% accuracy on the dev set (20% of the train set), highlighting the complexity of the task. The experimental setup for the data filtering classifier is described in Appendix A. Once trained, we apply the classifier to all LLM-generated data. After filtering, the dataset is reduced from 12,000 generated samples to 592 Sundanese and 628 Javanese high-quality samples. About half of the data generated by Claude and GPT-4o are retained, while a large portion of samples from other LLMs, especially Mixtral, are filtered out. More details on the filtered data proportions can be found in Appendix C.
Machine Translation
To compare LLM-assisted dataset construction with machine translation, we translate 1,000 instances from the original IndoCulture training set into 1,000 Javanese and 1,000 Sundanese examples. Both translations are generated using Google Translate [^3] https://translate.google.com/, accessed in September 2024., chosen for its strong translation performance in both languages. Since this dataset is intended for training, no human validation was conducted. From this point forward, we refer to it as MTtrain.
@@ 3.2 Test Set
Machine translation with human verification
As shown in Figure 1, we translate 500 randomly selected samples from the IndoCloze test set into Javanese and Sundanese using Google Translate. [^4] Google Translate was accessed in September 2024. Similar to the construction of training set, we choose Google Translate for its strong translation performance in both Javanese and Sundanese. To ensure quality, we employ two native speakers for each language and implement a two-stage quality control process. In Stage 1, the first worker manually corrects translation errors and localizes content by replacing entities (e.g., names, buildings, food) with culturally relevant alternatives. In Stage 2, a second worker validates the revised text and directly corrects any remaining errors from the first stage. This verification step ensures the accuracy and naturalness of the machine-translated test set. From this point forward, we refer to this data as MTtest.
Human-written Dataset
Each expert worker in Sundanese and Javanese is tasked with writing 600 short stories following the IndoCloze format: a four-sentence premise, a correct fifth sentence, and an incorrect fifth sentence. Stories are written based on 12 predefined topics, adhering to the same topic taxonomy used for training (see Figure 6 and Section 3.2). Each topic requires 50 stories, resulting in a total of 600 stories per language. [^5] See Appendix K for further details on the writing guidelines.
To ensure quality, each expert worker reviewed their peer’s written stories using the same pairing system as in the IndoCloze portion. The reviewing worker was presented with a premise and two randomized alternate endings from another worker’s story and was asked to identify the correct one. Instances incorrectly identified by the second worker were discarded, as they likely contained incorrect endings or exhibited ambiguity. After quality control, 529 Javanese and 595 Sundanese instances remained from the original 600 per language. From this point forward, we refer to this human-written dataset as HW.

### v2

# Culturally-Nuanced Story Generation for Reasoning in Low-Resource Languages: The Case of Javanese and Sundanese

@@ Abstract
Culturally grounded commonsense reasoning is underexplored in low-resource languages due to scarce data and costly native annotation. We test whether large language models (LLMs) can generate culturally nuanced narratives for such settings. Focusing on Javanese and Sundanese, we compare three data creation strategies: (1) LLM-assisted stories prompted with cultural cues, (2) machine translation from Indonesian benchmarks, and (3) native-written stories. Human evaluation finds LLM stories match natives on cultural fidelity but lag in coherence and correctness. We fine-tune models on each dataset and evaluate on a human-authored test set for classification and generation. LLM-generated data yields higher downstream performance than machine-translated and Indonesian human-authored training data. We release a high-quality benchmark of culturally grounded commonsense stories in Javanese and Sundanese to support future work.
Culturally-Nuanced Story Generation for Reasoning in Low-Resource Languages: The Case of Javanese and Sundanese
@@ 1 Introduction
Reasoning, the ability to draw conclusions, make inferences, and relate concepts, is a core evaluation target in recent LLM work (Dubey et al., 2024; OpenAI et al., 2024; Hurst et al., 2024; Almazrouei et al., 2023). Yet widely used English benchmarks such as StoryCloze (Mostafazadeh et al., 2016, 2017), WinoGrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019) encode Western norms. Because reasoning is culturally shaped, relying on machine-translated English datasets (Ponti et al., 2020; Lin et al., 2022) risks erasing local context.
Recent datasets for medium-resource languages (e.g., Indonesian Koto et al. (2024) and Arabic Sadallah et al. (2025)) add cultural grounding but mainly target sentence-level classification. Story-level commonsense, how people interpret events across narratives, remains underexplored in low-resource languages due to limited speaker access, high annotation costs, and scarce culturally relevant materials.
We address story comprehension in two underrepresented languages, Javanese and Sundanese, spoken by roughly 80M and 32M people respectively (Badan Pusat Statistik, 2025; Eberhard et al., 2025). Beyond sheer scale, both carry rich sociolinguistic systems: Sundanese encodes politeness and hierarchy phonologically, while Javanese employs elaborate speech levels (Wolff and Poedjosoedarmo, 1982). We adopt a StoryCloze-style setup (Mostafazadeh et al., 2016, 2017): given a four-sentence story, models either generate a plausible fifth sentence (generation) or choose the correct continuation from two options (classification).
We compare three dataset creation strategies for culturally grounded story comprehension: (1) LLM-assisted generation with culturally informed prompts, (2) machine translation from Indonesian benchmarks, and (3) native-authored stories. Each has distinct benefits, scalability, resource reuse, and authenticity, respectively. Human evaluation shows LLM outputs are culturally plausible and comparable to native-written stories.
Figure 1: Examples of human-written stories in Sundanese and Javanese. English translations in gray color are provided for reference. A cross (✗) indicates a culturally irrelevant ending, generated by LLMs
To assess cultural reasoning on narratives, we use the native-authored set as a zero-shot testbed. As in Figure 1, Indonesian/SEA-centric models, though covering the target languages, often produce culturally inappropriate endings, showing that language support alone is insufficient. We then fine-tune on synthetic data and find that LLM-generated training yields stronger results than machine-translated data for both classification and generation.
Our contributions are as follows:
  • We release the first benchmark for culturally grounded commonsense reasoning in Javanese and Sundanese, comprising 3.3K high-quality stories. This includes 1.12K human-written samples, 1K human-reviewed machine-translated texts, and 1.22K filtered LLM-generated samples.
  • We conduct extensive human evaluation of multiple dataset creation strategies, including (i) LLM-assisted generation, (ii) direct machine translation, (iii) culturally localized translation, and (iv) native-authored stories.
  • We evaluate model performance through zero-shot inference and supervised fine-tuning in both classification and generation settings to assess their cultural reasoning capabilities.
@@ 2 Related Works
2.1 Commonsense Reasoning in English Story Comprehension
Story comprehension in NLP involves reasoning over causal, temporal, and commonsense relations within narratives. The StoryCloze test, introduced by Mostafazadeh et al. (2016, 2017) is a landmark benchmark, requiring models to select the most plausible ending for a short four-sentence story. Many commonsense reasoning datasets, however, focus on sentence-level challenges include WinoGrande (Sakaguchi et al., 2021) for pronoun resolution, COPA (Gordon et al., 2012) for causal reasoning, and HellaSwag (Zellers et al., 2019) for adversarial sentence completion. While effective for probing localized reasoning, these do not capture broader discourse coherence or character motivations.
Recent work has shifted toward narrative-level reasoning with longer contexts and richer event dynamics. NarrativeQA (Kočiský et al., 2018) covers full books and movie scripts, CosmosQA (Huang et al., 2019) infers implicit causes and intentions, and TellMeWhy (Lal et al., 2021) targets causal and motivational “why” questions. Yet these remain English-centric and question-answering-oriented. Our work instead addresses narrative completion in low-resource languages, particularly in Javanese and Sundanese, providing a culturally grounded alternative to high-resource, English-dominant benchmarks.
2.2 Commonsense Reasoning in Languages Beyond English
Early multilingual commonsense benchmarks often extended English datasets via translation. XCOPA (Ponti et al., 2020) translated COPA into 11 typologically diverse languages, including Indonesian, while X-CSQA (Lin et al., 2021) adapted CommonSenseQA across languages. Although useful for cross-lingual evaluation, such resources inherit Anglocentric biases, as progress in English does not always transfer culturally or linguistically (Lin et al., 2022; Shwartz et al., 2020). Direct translations risk embedding English social contexts rather than local commonsense (Lin et al., 2021).
Story comprehension tasks like StoryCloze (Mostafazadeh et al., 2016) have been similarly extended. One such extenstion is XStoryCloze (Lin et al., 2022), by translating English narratives into multiple languages. Yet such approaches still struggle to capture culture-specific narrative norms.
For Indonesian, culturally grounded datasets such as COPAL-ID (Wibowo et al., 2024) and IndoCulture (Koto et al., 2024) model regional practices and norms across 11 provinces, advancing evaluation in a medium-resource language. However, they primarily focus on short-form, sentence-level reasoning such as multiple-choice or cloze-style questions, rather than full-narrative comprehension. Beyond Indonesia, CultureBank (Shi et al., 2024) compiles large-scale cultural knowledge from community narratives to support culturally aware language technologies, while CultureLLM (Li et al., 2025) incorporates cultural differences into LLMs via semantic data augmentation. However, these resources primarily focus on short-form, structured tasks rather than full-narrative comprehension. Our work fills this gap by introducing the first benchmark for story-level commonsense reasoning in low-resource languages, specifically Javanese and Sundanese, two of Indonesia’s most widely spoken local languages.
2.3 LLM-Generated Data Creation
One possible solution to tackle data scarcity in NLP is applying data augmentation (Feng et al., 2021; Ding et al., 2020; Ahmed and Buys, 2024; Liu et al., 2024; Yong et al., 2024; Guo and Chen, 2024; Liu et al., 2022), with LLMs increasingly used to produce high-quality synthetic data that complements or substitutes manual annotation. Most prior work targets classification tasks. For example, WANLI (Liu et al., 2022) used GPT-3 (Brown et al., 2020) to generate synthetic English natural language inference data (Bowman et al., 2015) refined by humans, while Yong et al. (2024) generated English sentiment and topic classification data before translating it into low-resource languages using bilingual lexicons.
For low-resource languages, Putri et al. (2024) employed GPT-4 (OpenAI et al., 2024) to create question-answering datasets, showing LLM potential in under-resourced settings. However, such efforts often overlook cultural reasoning and narrative coherence. Our work instead focuses on culturally nuanced story generation in Javanese and Sundanese, targeting story-level commonsense reasoning. We compare multiple data creation strategies, including LLM-assisted generation with open- and closed-weight models, machine translation from Indonesian, and native-authored stories.
@@ 3 Dataset Construction
Figure 2: Overall pipeline of dataset creation.
As in Figure 2, we build two parallel streams: training and test. For training, the IndoCloze (Koto et al., 2022) train split serves both as seeds for LLM-guided generation and as sources for machine translation into Javanese and Sundanese. For testing, we translate the IndoCloze test split and human-verify it for linguistic quality and cultural relevance, and we add a fully new set of native-authored stories from predefined topics. Each instance follows the StoryCloze format: a four-sentence premise with one correct and one incorrect ending.
Indonesian is chosen as the seed language for its national status and cultural proximity to Javanese and Sundanese. To ensure authenticity, native speakers authored and validated stories, embedding local names, places, foods, and customs. Following Mostafazadeh et al. (2016); Koto et al. (2022), the dataset targets everyday commonsense reasoning grounded in local culture
@@ 3.1 Training Data
We construct our training set using three strategies: (1) LLM-assisted data generation, (2) direct machine translation, and (3) machine translation followed by cultural localization using an LLM.
3.1.1 LLM-Assisted Data Generation
We synthesize training data with three open models: Gemma2-27B-it (Rivière et al., 2024), Llama3.1-70B (Dubey et al., 2024), Mixtral-8x7B-Instruct (Jiang et al., 2024), and three closed models: GPT-4o (Hurst et al., 2024; OpenAI et al., 2024), Cohere Command-R-Plus (Cohere, 2024), Claude-3-Opus (Anthropic, 2024).
For each language and LLM, we supply seed examples and topics. We manually translate 50 IndoCloze train samples into Javanese and Sundanese with cultural localization (e.g., foods like Gudeg and rituals like Sawéran). Topics are derived from IndoCulture (Koto et al., 2024). The full prompt appears in Figure 4.
To ensure that the generated examples align with the intended cultural context, we use cultural topics derived from IndoCulture (Koto et al., 2024). The full prompt used in this study is provided in Figure 4.
Each LLM produces 2,000 stories (1,000 per language), yielding six training sets. Per call, five random seeds guide generation at temperature 0.7; we repeat until each model reaches 1,000 valid samples per language ($\approx 166$ per topic on average).
To filter quality, we train an XLM-R binary classifier on 600 human-rated outputs (Section 4.2)11168.33% accuracy on the dev split (20% of train); setup in Appendix A.. Applied to all 12,000 generations, it retains 1.2K high-quality stories (592 Sundanese, 628 Javanese). Roughly half of Claude and GPT-4o outputs pass, while far fewer from other LLMs, especially Mixtral, survive (Appendix B).
3.1.2 Machine Translation (MTtrain)
We translate 1,000 IndoCulture training instances into Javanese and Sundanese (1,000 each) using Google Translate [^2] https://translate.google.com/, accessed September 2024. (see Appendix F for quality). As a training resource, this set is not human-validated. We refer to it as MTtrain.
3.1.3 Machine Translation + Localization (MTtrain+GPT4o)
We prompt GPT-4o to culturally localize the MT outputs, following the same guidelines given to human annotators. The model adapts names, events, foods, settings, and social norms to Javanese/Sundanese contexts. This probes whether LLMs can align literal translations with local cultural values to produce more authentic, context-appropriate narratives.
@@ 3.2 Test Set
3.2.1 Machine translation with human verification (MTtest+Human)
As shown in Figure 2, we translate 500 randomly selected samples from the IndoCloze test set into Javanese and Sundanese using Google Translate. [^3] Google Translate was accessed in September 2024 To ensure accuracy and naturalness of the machine-translation, we employ two native speakers for each language and implement a two-stage quality control process. In Stage 1, the first worker manually corrects translation errors and localizes content by replacing entities (e.g., names, buildings, food) with culturally relevant alternatives. [^4] Note that while we applied cultural localization, not all examples could be fully adapted to Javanese or Sundanese contexts, as some stories reflect general Indonesian cultural elements that are not specific to either group. In Stage 2, a second worker validates the revised text and directly corrects any remaining errors from the first stage. From this point forward, we refer to this data as MTtest+Human.
3.2.2 Human-written Dataset (HW)
Each expert worker in Sundanese and Javanese is tasked with writing 600 short stories following the IndoCloze format: a four-sentence premise, a correct fifth sentence, and an incorrect fifth sentence. Stories are written based on 12 predefined topics, adhering to the same topic taxonomy used for training (see Section 3.2).See Appendix E for further details on the writing guidelines.
To ensure quality, each expert worker reviewed their peer’s written stories. The reviewing worker was presented with a premise and two randomized alternate endings from another worker’s story and was asked to identify the correct one. Instances incorrectly identified by the second worker were discarded, as they likely contained incorrect endings or exhibited ambiguity. After quality control, 529 Javanese and 595 Sundanese instances remained from the original 600 per language. From this point forward, we refer to this human-written dataset as HW.